# A Multi-Fidelity Parametric Framework for Reduced-Order Modeling using Optimal Transport-based Interpolation: Applications to Diffused-Interface Two-Phase Flows

### 04 Mar 2026 v1

@@ Abstract.
This work introduces a data-driven, non-intrusive reduced-order modeling (ROM) framework that leverages Optimal Transport (OT) for multi-fidelity and parametric problems. Building upon the success of displacement interpolation for data augmentation in handling nonlinear dynamics, we extend its application to more complex and practical scenarios. The framework is designed to correct a computationally inexpensive low-fidelity (LF) model to match an accurate high-fidelity (HF) one by capturing its temporal evolution via displacement interpolation while preserving the problem’s physical consistency. The framework is further extended to address systems dependent on a physical parameter, for which we construct a surrogate model using a hierarchical, two-level interpolation strategy. First, it creates synthetic HF checkpoints via displacement interpolation in the parameter space. Second, the residual between these synthetic HF checkpoints and a true LF solution is interpolated in the time domain using the multi-fidelity OT-based methodology. This strategy provides a robust and efficient way to explore the parameter space and to obtain a refined description of the dynamical system. The potential of the method is discussed in the context of complex and computationally expensive diffuse interface methods for two-phase flow simulations, which are characterized by moving interfaces and nonlinear evolution, and challenging to be dealt with traditional ROM techniques.
@@ 1. Introduction
Many engineering and scientific phenomena of practical interest can be accurately described by mathematical models based on partial differential equations (PDEs). The numerical solution of such PDEs is generally performed through spatio-temporal discretization, using methods like the Finite Element (FE) or Finite Volume (FV) methods, which transform the continuous problem into a system with a finite number $N_{h}$ of degrees of freedom [57]. An accurate discretization of general PDEs often yields high-dimensional solution vectors, meaning that a proper description of their behavior requires a computationally intractable number of degrees of freedom ($N_{h}\gg 1$). For time-dependent problems, further discretization in time results in a sequence of large algebraic systems to be solved, compounding the computational burden.
@@
In our previous work [38], we introduced a ROM framework that addresses these challenges by incorporating the concept of Displacement Interpolation (DI) from Optimal Transport (OT) theory [73, 52, 64] to enhance the representation of nonlinear dynamics in complex systems. OT, originally developed by Monge and Kantorovich, concerns the matching and transport of probability distributions by identifying the most cost-effective way to transform one into another. Unlike classical $L^{2}$-based distance metrics that compare distributions pointwise, OT considers the “work” required to move a source mass distribution to match a target distribution, inherently accounting for the underlying geometry of the space. A fundamental concept in OT is the Wasserstein distance, which quantifies the minimum cost of transporting one distribution to another [21, 7, 20]. This distance defines a metric in the space of probability measures, endowing it with a Riemannian-like structure that allows for geometric operations such as the definition of geodesics and barycenters [1, 42]. The geometric perspective provided by OT is particularly attractive for ROMs concerned with the treatment of features that translate or deform, as the Wasserstein distance can capture these transformations in a more natural way compared to traditional Euclidean-based metrics. By interpreting solution snapshots as probability distributions and leveraging displacement interpolation—the construction of geodesic paths in Wasserstein space—our method generates physically plausible synthetic snapshots between sparse high-fidelity checkpoints, addressing the challenge of data scarcity while enabling continuous-time prediction [33, 18, 38].
In this paper, we extend this OT-based methodology to address two challenges in computational modeling that arise frequently in practice, namely the multi-fidelity and the parametric modeling.
@@
Indeed, in the first context, the goal is to correct computationally inexpensive low-fidelity (LF) approximations with limited accuracy, for instance due to a coarser mesh or simplified physics, using information from a few expensive high-fidelity (HF) simulations [49, 30, 17]. Here, we propose a new approach where this information is retrieved by interpolating in time the residual field between HF and LF solutions via OT, leveraging the geometric structure of Wasserstein space to track the movement of discrepancy features.
@@
Moreover, many physical systems depend on a set of parameters $\bm{\mu}$, such as material properties or boundary conditions, and the goal is to build a ROM capable of rapid prediction for any $\bm{\mu}$ within a given parameter domain. To investigate efficiently different system’s configurations, we introduce a hierarchical, two-level interpolation strategy to simultaneously deal with both parameter and time dependency. In the first level, we use the OT-based displacement interpolation across the parameter space to generate synthetic HF snapshots for unseen parameter value from HF checkpoints computed at nearby training points. In the second level, we compute the residual between these physically consistent HF synthetic data and true LF simulations at the test parameter, and then apply our multi-fidelity framework to interpolate this residual in time. This approach avoids the construction of a monolithic ROM over the joint time-parameter space, which can be prohibitively complex. These two extensions, the Multi-Fidelity OT-ROM (MF-OT-ROM) and the Parametric Multi-Fidelity OT-ROM (PMF-OT-ROM), broaden the applicability of our original method.
We frame our discussion within the context of two-phase flow modelling using the conservative Allen-Cahn equation [14]. More broadly, diffuse-interface methods have gained significant attention in the computational fluid dynamics community due to their versatility across different numerical schemes, governing equations, and flow regimes (incompressible or compressible, laminar or turbulent), as well as their relative ease of implementation [65]. The widespread adoption of these approaches has enabled the simulation of highly complex two-phase flows, which are often characterized by strongly coupled phenomena, including interfacial mass and heat transfer [44], phase change [63], boiling [61, 32, 74], turbulence [68, 60], and strong compressibility effects [31, 71, 16], among others. These capabilities make diffuse-interface methods particularly suitable for tackling multi-physics scenarios that are challenging to address with sharp-interface models.
In particular, recent developments of the conservative Allen-Cahn equation and its variants have yielded numerical methods with several desirable properties, including enhanced mass conservation [69], strict boundedness [43], and efficient interface sharpening [2, 35]. Despite the growing literature on high-fidelity simulations employing these methodologies, the integration of reduced-order modelling strategies remains largely unexplored in this field. Two-phase flows fall into the broader category of advection-dominated problems, for which classical linear ROMs often struggle to maintain accuracy and robustness, especially in the presence of complex interfacial dynamics and multi-scale interactions.
In this work, we extend the methodological developments of OT-based ROMs and investigate their applicability to two-phase flow simulations. By doing so, we aim to bridge the gap between high-fidelity diffuse-interface models and computationally efficient ROMs, providing new avenues for reduced-order modelling in a challenging and highly relevant class of multi-physics problems.
@@
The remainder of the paper is structured as follows. Section 2 provides the necessary mathematical background on reduced-order modeling and optimal transport theory, and summarizes the core OT-ROM methodology from our previous work [38], which serves as the foundation for the extensions presented here. Section 3 introduces multi-fidelity residual interpolation, while Section 4 presents parametric displacement interpolation, applicable to both single- and multi-fidelity settings. Section 5 introduces the conservative Allen-Cahn equation coupled within the five-equation model, while Section 6 is reserved for numerical experiments. Finally, Section 7 provides concluding remarks and directions for future research.
2. Preliminaries
@@ 2.1. Reduced-Order Modeling: The Offline-Online Paradigm
In particular, dimensionality reduction techniques aim to identify the underlying low-dimensional structure of the dominant pattern of the solution manifold characterized by the evolution in time and the variation w.r.t. input parameters. Toward this goal, the system dynamics is projected onto this reduced set of coordinates, building a surrogate model that is much faster to solve than its HF counterpart, and retain the essential physical characteristics and accuracy.
@@
where $\epsilon$ is a small tolerance. The rate at which the singular values decay determines how many modes are needed for an accurate approximation, and thus the reliability of the ROM ansatz. As described by the Kolmogorov n-width decay analysis [15, 48, 51], for advection-dominated problems with moving features this decay is typically slow, necessitating a large number of modes and diminishing the computational advantage of the ROM.
@@ 2.2. Optimal Transport and Displacement Interpolation
Optimal Transport (OT) theory provides a geometric framework for comparing and morphing probability distributions [73, 52]. At its core, OT addresses the problem of finding the most efficient way to transform one distribution of mass into another, where efficiency is measured by a total transportation cost. The theory was originally posed by Monge, looking for the cheapest way to transport soil to fill a hole [45], and later relaxed by Kantorovich [37] allowing mass from a single source point to be split and distributed among multiple destinations, leading to the modern theory of finding transport maps, for a given cost, which are optimal in matching source and target distributions. Figure 1 illustrates the difference between the two formulations in terms of mass splitting and the resulting transport plans.
@@
We specialize directly to the discrete setting, which is of practical importance for numerical applications. Let the source and target measures be discrete probability measures with finite supports, written as weighted sums of Dirac masses:
@@
$C(\mu,\nu)=\min_{\bm{\pi}\in\Pi(\mu,\nu)}\sum_{i=1}^{n}\sum_{j=1}^{m}C_{ij}\pi_{ij}=\min_{\bm{\pi}\in\Pi(\mu,\nu)}\langle\mathbf{C},\bm{\pi}\rangle_{F},$ (7)
@@
$W_{p}(\mu,\nu)=\left(\min_{\bm{\pi}\in\Pi(\mu,\nu)}\sum_{i,j}\|{\bm{x}}_{i}-{\bm{y}}_{j}\|^{p}\,\pi_{ij}\right)^{1/p}.$ (8)
The space of probability measures equipped with this metric inherits a rich geometric structure. Unlike classical $L^{2}$-based metrics, which only measure pointwise differences, the Wasserstein distance is sensitive to the spatial displacement of features. Two distributions representing a feature and its translated counterpart are close in Wasserstein distance, reflecting the physical intuition that small translations should correspond to small distances. Then, it is clear that the choice of the underlying metric is particularly relevant in the context of physical fields exhibiting sharp contrasts, discontinuous features and small compact supports.
A fundamental property of spaces endowed with the Wasserstein metric is that they admit geodesics, i.e., shortest paths between measures. Given an optimal transport plan $\bm{\pi}^{*}$ between $\mu_{0}$ and $\mu_{1}$, one can construct a geodesic path $\{\mu_{t}\}_{t\in[0,1]}$ called the displacement interpolation or McCann interpolation:
@@
The displacement interpolation provides a geometrically natural way to “morph” between distributions, and given it nice properties in dealing with moving features, we leverage this as the key tool for generating synthetic snapshots describing the time- and parametric-evolution of physical fields.
@@
where the Gibbs kernel is defined as:
@@
The concept also generalizes to multiple measures via Wasserstein barycenters. Given measures $\{\mu_{k}\}_{k=1}^{K}$ with weights $\{\lambda_{k}\}$ summing to 1, the barycenter $\bar{\mu}$ minimizes:
$\bar{\mu}=\operatorname*{arg\,min}_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\sum_{k=1}^{K}\lambda_{k}W_{2}^{2}(\mu,\mu_{k}).$ (17)
This provides a geometrically meaningful “average” of distributions and can be used for multi-source parametric interpolation when more than two training parameters are involved.
@@ 2.3. OT-Based Displacement Interpolation for ROM
In this final section, we summarize the core OT-ROM methodology introduced in our previous work [38], which serves as the foundation for the extensions presented in this paper. The central idea is to use displacement interpolation to generate synthetic snapshots between sparse high-fidelity checkpoints, addressing data scarcity in ROM construction.
Consider a time-dependent problem whose solution ${\bm{u}}(t)$ is available at a sparse set of $N_{c}$ checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$. The goal is to construct a continuous-in-time approximation ${\bm{u}}_{\text{synth}}(t)$ for any $t\in[t_{1},t_{N_{c}}]$ via the displacement interpolation framework. The procedure, allowing for the efficient evaluation of time-trajectory snapshots and data-augmentation tasks, is structured as follows:
@@
This approach exploits the geometric structure of Wasserstein space to transport features between their positions at time $t_{k}$ and $t_{k+1}$, rather than simply blending them pointwise as in linear interpolation, which would be completely ineffective in the case of moving features with compact support. Moreover, the resulting synthetic snapshots preserve mass and follow geodesic paths in the space of probability measures, and thus, despite being a data-driven approach, this promotes physical and geometric consistency with the underlying PDEs.
@@
Instead of choosing the standard linear time mapping for the virtual time $\alpha$, one could also consider the more accurate nonlinear time mapping introduced in [38], which minimizes the $L^{2}$ error between synthetic and true snapshots. In general, accounting for multi-scale behavior in time could prove crucial for the accuracy of the online reconstruction, but for the problems considered here the linear mapping has been sufficient to provide accurate results.

### 27 Mar 2026 v2

# A Multi-Fidelity and Parametric Reduced-Order Modeling Framework with Optimal Transport-based Interpolation: Applications to Diffused-Interface Two-Phase Flows

@@ Abstract.
This work introduces a data-driven, non-intrusive reduced-order modeling (ROM) framework that leverages Optimal Transport (OT) for multi-fidelity and parametric problems in two-phase flows modelling. Building upon the success of displacement interpolation for data augmentation in handling nonlinear dynamics, we extend its application to more complex and practical scenarios. The framework is designed to correct a computationally inexpensive low-fidelity (LF) model to match an accurate high-fidelity (HF) one by capturing its temporal evolution via displacement interpolation while preserving the problem’s physical consistency. The framework is further extended to address systems dependent on a physical parameter, for which we construct a surrogate model using a hierarchical, two-level interpolation strategy. First, it creates synthetic HF checkpoints via displacement interpolation in the parameter space. Second, the residual between these synthetic HF checkpoints and a true LF solution is interpolated in the time domain using the multi-fidelity OT-based methodology. This strategy provides a robust and efficient way to explore the parameter space and to obtain a refined description of the dynamical system. The potential of the method is discussed in the context of complex and computationally expensive diffuse-interface methods for two-phase flow simulations, which are characterized by moving interfaces and nonlinear evolution, and challenging to be dealt with traditional ROM techniques.
@@ 1. Introduction
Many engineering and physical phenomena of practical interest can be accurately described by mathematical models based on partial differential equations (PDEs). The numerical solution of such PDEs is generally performed through spatio-temporal discretization, using methods like the Finite Element (FE) or Finite Volume (FV) methods, which transform the continuous problem into a system with a finite number $N_{h}$ of degrees of freedom [71]. An accurate discretization of general PDEs often yields high-dimensional solution vectors, meaning that a proper description of their behavior requires a computationally intractable number of degrees of freedom ($N_{h}\gg 1$). For time-dependent problems, further discretization in time results in a sequence of large algebraic systems to be solved, compounding the computational burden.
@@
In our previous work [48], we introduced a ROM framework that addresses these challenges by incorporating the concept of Displacement Interpolation (DI) from Optimal Transport (OT) theory [88, 65, 78] to enhance the representation of nonlinear dynamics in complex systems. OT, originally developed by Monge [57] and Kantorovich [46], concerns the matching and transport of probability distributions by identifying the most cost-effective way to transform one into another. Unlike classical $L^{2}$-based euclidean distance metrics that compare distributions pointwise, OT considers the “work” required to move a source mass distribution to match a target distribution, inherently accounting for the underlying geometry of the space.
The fundamental concept behind this is the so-called Wasserstein distance, which quantifies the minimum cost of transporting one distribution to another [24, 8, 23]. This distance defines a metric in the space of probability measures, endowing it with a Riemannian-like structure that allows for geometric operations such as the definition of geodesics and barycenters [1, 54]. The geometric perspective provided by OT is particularly attractive for ROMs concerned with the treatment of features that translate or deform, as the Wasserstein distance can capture these transformations in a more natural way compared to traditional Euclidean-based metrics. By interpreting solution snapshots as probability distributions and leveraging displacement interpolation—the construction of geodesic paths in Wasserstein space—our method generates physically plausible synthetic snapshots between sparse high-fidelity checkpoints, addressing the challenge of data scarcity while enabling continuous-time prediction [40, 20, 48].
@@
In this paper, we extend this OT-based methodology to address two challenges in computational modeling that arise frequently in practice, namely the multi-fidelity and the parametric modeling. Indeed, in the first context, the goal is to correct computationally inexpensive low-fidelity (LF) approximations with limited accuracy, for instance due to a coarser mesh or simplified physics, using information from a few expensive high-fidelity (HF) simulations [62, 35, 19]. Here, we propose a new approach where this information is retrieved by interpolating in time the residual field between HF and LF solutions via OT, leveraging the geometric structure of Wasserstein space to track the movement of discrepancy features.
@@
Moreover, many physical systems depend on a set of parameters $\bm{\mu}$, such as fluids and material properties or boundary/initial conditions, and the goal is to build a ROM capable of rapid prediction for any $\bm{\mu}$ within a given parameter domain. To investigate efficiently different system’s configurations, we introduce a hierarchical, two-level interpolation strategy to simultaneously deal with both parameter and time dependency. In the first level, we use the OT-based displacement interpolation across the parameter space to generate synthetic HF snapshots for unseen parameter value from HF checkpoints computed at nearby training points. In the second level, we compute the residual between these physically consistent HF synthetic data and true LF simulations at the test parameter, and then apply our multi-fidelity framework to interpolate this residual in time. This approach avoids the construction of a monolithic ROM over the joint time-parameter space, which can be prohibitively complex. These two extensions, the Multi-Fidelity OT-ROM (MF-OT-ROM) and the Parametric Multi-Fidelity OT-ROM (PMF-OT-ROM), broaden the applicability of our original method.
@@
The remainder of the paper is structured as follows. Section 2 presents a brief review of diffuse-interface methods for two-phase flows, with a particular focus on surrogate modelling, serving as the application context for the proposed methodologies. Section 3 provides the necessary mathematical background on reduced-order modeling and optimal transport theory. Section 3.3 summarizes the core OT-ROM methodology from our previous work [48], which serves as the foundation for the extensions presented here. Section 4 introduces multi-fidelity residual interpolation, while Section 5 presents parametric displacement interpolation, applicable to both single- and multi-fidelity settings. Section 6 introduces the conservative Allen-Cahn equation coupled within the five-equation model, while Section 7 is reserved for numerical experiments. Finally, Section 8 provides concluding remarks and directions for future research.
2. Diffuse-interface two-phase modelling
We frame our discussion within the context of two-phase flow modelling using the conservative Allen-Cahn equation [16] coupled with the five-equation model [3]. More broadly, diffuse-interface methods have gained significant attention in the computational fluid dynamics community due to their versatility across different numerical schemes, mesh anisotropies, governing equations and flow regimes (from incompressible to compressible, from laminar to turbulent), as well as their relative ease of implementation (see [79] for an extensive review). In these approaches, the interface between the two fluids is represented as a thin but finite transition layer described by a continuous phase-field variable, thereby avoiding the need for explicit interface tracking or reconstruction as it custom in Volume-of-Fluid (VOF) or Level-Set (LS) methods. This modelling paradigm offers significant advantages in terms of numerical robustness and flexibility when dealing with complex interfacial topological changes such as breakup, coalescence, or strong deformation.
The widespread adoption of diffuse-interface methods has enabled the simulation of highly complex two-phase flows, which are often characterized by strongly coupled multi-physics phenomena. These include interfacial mass and heat transfer [56], phase change [30, 38, 77], boiling processes [75, 39, 89], turbulent interfacial dynamics [82, 74], and strong compressibility effects [36, 18, 90]. The ability to address such tightly coupled physical processes within a unified modelling framework makes diffuse-interface methods particularly attractive for the simulation of realistic engineering and scientific applications. At the same time, the increasing fidelity and resolution of these simulations often come with a substantial computational cost, motivating the development of efficient surrogate and reduced-order modelling strategies.
In particular, recent developments of the conservative Allen-Cahn equation and its variants have led to numerical methods exhibiting several desirable properties, including improved mass conservation [83], strict boundedness [55, 69, 37], and effective interface sharpening [2, 43, 85]. These advancements have contributed to the growing popularity of formulations based on the conservative Allen-Cahn equations for high-fidelity two-phase flow simulations. Despite the growing body of literature on high-fidelity simulations based on these approaches, the integration of reduced-order modelling techniques in this context has only recently begun to receive attention [91, 45, 29, 21]. Developing robust and accurate ROMs for such systems remains particularly challenging due to the strongly nonlinear and transport-dominated nature of the underlying dynamics, such as the case of two-phase flows being characterized by complex interfacial dynamics, sharp gradients, and multi-scale interactions. In this setting, we aim at showing the potentiality of OT-ROM techniques to efficiently capture transport-dominated phenomena.
Indeed, in the context of diffuse-interface models, several properties of the phase-field variable make OT particularly well suited for this class of problems. The phase-field variable distinguishing the two phases is in fact naturally bounded between zero and one. Moreover, in incompressible or low-Mach-number compressible flows, the total volume associated with the phase-field is preserved over time. These properties resemble those of probability distributions within the OT formalism, where quantities are non-negative and globally conserved. This analogy provides a natural connection between diffuse-interface representations of two-phase flows and OT-based reduction strategies, further confirming that OT-based techniques may offer a promising pathway for developing robust and efficient reduced-order models in this setting.
Finally, in this work we extend the methodological developments of OT-based ROMs and investigate their applicability to two-phase flow simulations. By doing so, we aim to bridge the gap between high-fidelity diffuse-interface models and computationally efficient ROMs. The resulting framework leverages OT to construct reduced representations that remain accurate even in the presence of strongly transport-dominated dynamics and evolving interfaces.
3. OT-ROMs Preliminaries
@@ 3.1. Reduced-Order Modeling: The Offline-Online Paradigm
In particular, dimensionality reduction techniques aim to identify the underlying low-dimensional structure of the dominant pattern of the solution manifold characterized by the evolution in time and the variation w.r.t. input parameters. Toward this goal, the system dynamics is projected onto a reduced set of coordinates, building a surrogate model that is much faster to solve than its HF counterpart, and retains the essential physical characteristics and accuracy.
@@
where $\epsilon$ is a small tolerance. The rate at which the singular values decay determines how many modes are needed for an accurate approximation, and thus the reliability of the ROM ansatz. As described by the Kolmogorov n-width decay analysis [17, 61, 64], for advection-dominated problems with moving features this decay is typically slow, necessitating a large number of modes and thus compromising the computational advantage of the ROM.
@@ 3.2. Optimal Transport and Displacement Interpolation
Optimal Transport (OT) theory provides a geometric framework for comparing and morphing probability distributions [88, 65]. At its core, OT addresses the problem of finding the most efficient way to transform one distribution of mass into another, where efficiency is measured by a total transportation cost. The theory was originally posed by Monge [57], looking for the cheapest way to transport soil to fill a hole, and later relaxed by Kantorovich [46] allowing mass from a single source point to be split and distributed among multiple destinations. This latter interpretation led to the modern theory of finding transport maps, for a given cost, which are optimal in matching source $\mu$ and target $\nu$ distributions. Figure 1 illustrates the difference between the two formulations in terms of mass splitting and the resulting plans $\bm{\pi}$ to transport the two distributions.
@@
Despite more general formulations are available [9, 52, 49], here we specialize our discussion directly to the discrete setting, which is of practical importance for numerical applications. Let the source and target measures be discrete probability measures with finite supports, written as weighted sums of Dirac masses $\delta$:
@@
$C(\mu,\nu)\doteq\min_{\bm{\pi}\in\Pi(\mu,\nu)}\sum_{i=1}^{n}\sum_{j=1}^{m}C_{ij}\pi_{ij}=\min_{\bm{\pi}\in\Pi(\mu,\nu)}\langle\mathbf{C},\bm{\pi}\rangle_{F},$ (7)
@@
$W_{p}(\mu,\nu)\doteq\left(\min_{\bm{\pi}\in\Pi(\mu,\nu)}\sum_{i,j}\|{\bm{x}}_{i}-{\bm{y}}_{j}\|^{p}\,\pi_{ij}\right)^{1/p}.$ (8)
The space of probability measures equipped with this metric inherits a rich geometric structure. Unlike classical $L^{2}$-based metrics, which only measure pointwise differences, the Wasserstein distance is sensitive to the spatial displacement of features. Indeed, two distributions representing a feature and its translated counterpart are close in Wasserstein distance, reflecting the physical intuition that small translations should correspond to small distances, but very far in the standard Euclidean metric. Then, it is clear that the choice of the underlying metric is particularly relevant in the context of physical fields exhibiting sharp contrasts, discontinuous features and small compact supports.
A fundamental property of spaces endowed with the Wasserstein metric is that they admit geodesics, i.e., shortest paths between measures. Given an optimal transport plan $\bm{\pi}^{*}$ between $\mu_{0}$ and $\mu_{1}$, one can construct a geodesic path $\{\mu_{t}\}_{t\in[0,1]}$ called the displacement interpolation or McCann interpolation [54, 40, 48]:
@@
The displacement interpolation provides a geometrically natural way to “morph” between distributions, and given its nice properties in dealing with moving features, we leverage this as the key tool for generating synthetic snapshots describing the time- and parametric-evolution of physical fields.
@@
where $\mathbf{K}$ is the Gibbs kernel defined as:
@@ 3.3. OT-Based Displacement Interpolation for ROM
Finally, we summarize the core OT-ROM methodology introduced in our previous work [48], which serves as the foundation for the extensions presented in this paper. The central idea is to use displacement interpolation to generate synthetic snapshots between sparse high-fidelity checkpoints, addressing data scarcity in ROM construction.
Consider a time-dependent problem whose solution ${\bm{u}}(t)$ is available at a sparse set of $N_{c}$ checkpoint times $\{t_{k}\}_{k=1}^{N_{c}}$, with $N_{c}\ll N_{T}$. The goal is to construct a continuous-in-time approximation ${\bm{u}}_{\text{synth}}(t)$ for any $t\in[t_{1},t_{N_{c}}]$ via the displacement interpolation framework. The procedure, allowing for the efficient evaluation of time-trajectory snapshots and data-augmentation tasks, is structured as follows:
@@
This approach exploits the geometric structure of Wasserstein space to transport features between their positions at time $t_{k}$ and $t_{k+1}$, rather than simply blending them pointwise as in linear interpolation, which would be completely ineffective in the case of moving features with compact support. Moreover, the resulting synthetic snapshots preserve mass and follow geodesic paths in the space of probability measures, and thus, despite being a data-driven approach, this promotes physical and geometric consistency with the underlying PDE.
@@
Instead of choosing the standard linear time mapping for the virtual time $\alpha$, one could also consider the more accurate nonlinear time mapping introduced in [48], which minimizes the $L^{2}$ error between synthetic and available true snapshots. In general, accounting for multi-scale behavior in time could prove crucial for the accuracy of the online reconstruction, but for the problems considered here the linear mapping has been sufficient to provide accurate results.